OpenStudy (anonymous):

Let log2=0.428 and log3=0.519, what is the approximate value of log24?

OpenStudy (anonymous):

log 24 =log(2)(2)(2)(3) = 3log2 + log 3 = 3(0.428)+0.519

OpenStudy (campbell_st):

24 = 2^3 x 3 \[\log(2^3 \times3) = \log 2^3 + \log 3 = 3 \times \log(2) + \log 3\]

OpenStudy (anonymous):

thank yalll